OpenStudy (anonymous):

Evaluate(x)= 4x-1 for=3

OpenStudy (anonymous):

y = 4(–1) – 3 = –4 – 3 = –7

OpenStudy (anonymous):

but double check bby

OpenStudy (pawanyadav):

Do you mean 4x-1=3 4x=4 so x=1

OpenStudy (anonymous):

oh yea